A.     List all the subsets of the following sets.

2. {1,2,Ø}

8. {{0,1},{0,1,{2}},{0}}


B. Write out the following sets by listing their elements between braces.

12. {X:X ⊆{3,2,a} and │X│=1


C. Decide if the following statement is tru or false. Explain.

14. R2⊆R3

1
Expert's answer
2021-03-22T11:36:26-0400

A.     Let us list all the subsets of the following sets.


2. The subsets of the set "\{1,2,\emptyset\}" are "\emptyset, \{0\},\{1\},\{\emptyset\},\{1,2\},\{1,\emptyset\},\{2,\emptyset\}, \{1,2,\emptyset\}."


8. The subsets of the set "\{\{0,1\},\{0,1,\{2\}\},\{0\}\}" are "\emptyset, \ \{\{0,1\}\},\ \ \{\{0,1,\{2\}\}\},\ \{\{0\}\}, \ \{\{0,1\},\{0,1,\{2\}\}\},\newline \{\{0,1\},\{0\}\},\ \{\{0,1,\{2\}\},\{0\}\}, \ \{\{0,1\},\{0,1,\{2\}\},\{0\}\}."


B. Let us write out the following sets by listing their elements between braces.


12. "\{X:X \subset \{3,2,a\}\text{ and }│X│=1\}=\{\{3\},\{2\},\{a\}\}."


C. Let us decide if the following statement is true or false.


14. The statement "\mathbb R^2⊆ \mathbb R^3" is false because the elements of "\mathbb R^2" are pairs "(x,y)" of real numbers "x" and "y", but the the elements of "\mathbb R^3" are triples "(x,y,z)" of real numbers "x,y" and "z."